OpenStudy (anonymous):

Can someone give me example expression i can work out? EXP-there are 9 dougnuts in a box, 2 of them are choccolate, give me an expressiong of how many doughnuts are not chocolate.

OpenStudy (anonymous):

lot's of misspelling sorry, just ignore it

OpenStudy (anonymous):

how about \(9-2=7\) ?

OpenStudy (anonymous):

you have to use a variable in there.

OpenStudy (anonymous):

where the 9 is?

OpenStudy (anonymous):

or where the answer is?

OpenStudy (anonymous):

guess you could say \(x=9-2\) if you like kind of weird though because you are acting as if you don't know what \(9-2\) gives you

OpenStudy (anonymous):

7 duh

OpenStudy (anonymous):

i'm not a retard u know

OpenStudy (anonymous):

is the variable supposed to be the answer? how many are not chocolate? it can't be the number of donuts because you are told there are 9 of them and 9 is not a variable i guess i don't understand what the question is actually asking

OpenStudy (anonymous):

i would go with \(x=9-2\)

OpenStudy (anonymous):

where does this question come from? on line class?

OpenStudy (anonymous):

no my dad

OpenStudy (anonymous):

i didn't mean "you" don't know what \(9-2\) is i meant the question as acting AS IF you didn't know what \(9-2\) is

OpenStudy (anonymous):

ask him what the variable is supposed to represent the number of donuts the number of chocolate donuts the number of not chocolate donuts then maybe gently suggest that there are no variables in this problem just the numbers 9, 2 and 7

OpenStudy (anonymous):

i think it's 9-c c=2, and woudn't that give u the amount of dougnuts that's not chocolate??

OpenStudy (anonymous):

ok